Solve.
A high school basketball court is 84 feet long and 50 feet wide. What is the diagonal length of the court?

Answer by Jstrasner(112)

Hey,
For this one we use the pythagorean theorem:
a^2 + b^2 = c^2
Therefore:
84^2 + 50^2 = c^2
7056 + 2500 = c^2
9556 = c^2
97.75 = c
Therefore:
The diagonal is about 98 feet long
I hope this helps!